40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 47 of 52  Not logged in ELOG logo
ID Date Author Type Category Subject
  293   Wed Aug 18 20:04:53 2010 FrankSummaryDAQVME 3123 card NOT broken

it turned out that the ADC card is not broken. Instead it's a common-mode range problem of the inputs.
As we are using several PDs on the table powered by individual (non-grounded) power supplies (Thorlabs photodiodes), these signals are floating around.
And turns out that for some reason since a couple of days the common-mode voltage was sometimes larger than the input can handle (+/-12V) and so strange things happen to the signals from time to time.
In order to fix that i added 1MOhm resistors from the negative input to analog ground of the adc-card to prevent that input from floating around. The negative input is usually ground as all our signal are single-ended referred-to-ground signals; there are no differential transmitted signals for those devices) and so this should work fine. I only added one resistor per device connected to the DAQ, meaning the temp box with four signals has only one (the first one) connected to ground as all other ones have the same ground and so they can't floaf different anymore.

That fixed all problems...

  292   Wed Aug 18 16:33:52 2010 FrankSummaryDAQ3123 card broken ?

the 3123 card (16bit input) seems to be broken. That's the card which also samples the PMC transmitted light which fluctuates periodically since a couple of days...
I changed the database to read the temp sensors for both cavities and when testing the inputs i figured out that it's reading only 8.97V of the 10 i put in, same for other values...
soft and hard resetting the crate doesn't change anything...

gonna replace the card...

  291   Wed Aug 18 12:19:17 2010 FrankPhotos beat setup panorama picture incl beams

i know the the asymmetry is not nice but we made the first beat setup symmetric and suffered a lot from the (former) little space behind the cavities.
Also one of the cavities is tilted and so the beam height is about .5in different.
Now as me moved both cavities a bit further away from the end of the table we should re-think about changing it to a symmetric setup. I think it's a good idea.
Anyway, shouldn't we just see all the resonances of the mounts in the spectrum, not a broadband noise hump? or do you expect lots of scatter from all the unnecessary optics?

you are right with the power, i reduced it by a factor of 10 and 100 but no change in noise spectrum (almost absolutely the same). But if i reduce the power some pickup of the 35.5MHz shows up in the signal (which is tiny), which i didn't see before. But it's ground loop related as you can change it by connecting/disconnecting some more channels to the scope e.g.. About the gain: i didn't measure the TIA gain of the PD, but the manual says 24e3 in the text, and 40e3 in the table at the end. Which one is right?

Another problem is that the bandwidth is only 125MHz, but we are looking at 160MHz. That shouldn't be much of a problem, only the noise of the PD increases and we are more sensitive to changes in the TF of the PD due to almost anything like power fluctuations, power supply fluctuations, temperature etc... So that probably the best argument right now to add the second AOM to reduce the beat frequency to something we can handle much better

Quote:

The panorama is a good idea. We should make it a mandatory step whenever we make any change to the setup.

The asymmetry in the paths of the transmitted beams is dis-satisfying though. I would have tried to just take the transmitted beam and interfere them via short path lengths and no lenses. The cavity eigenmodes of each cavity should already be matched to ~1%. If the path length from the cavities to the BS can be made equal, the overlap should also be good. In general, it is hard to make a low phase noise setup with a long path length.

To ensure low backscatter from the transmission RFPDs back into the cavity, the beat signal PD should be placed slightly ahead of the focus of the final lens.

And since the Z=40 kOhms for the 1811, the light power from each cavity should be made ~200 mV / 40000 ~ 5 uW. I know its tiny, but otherwise the 1811 is not going to be very linear. And the RF signal going into a level 13 mixer ought to be less than 300 mVrms.

The better option is to use a RFPD with only ~1 kOhm of RF transimpedance...

 

  290   Wed Aug 18 03:15:20 2010 ranaPhotos beat setup panorama picture incl beams

The panorama is a good idea. We should make it a mandatory step whenever we make any change to the setup.

The asymmetry in the paths of the transmitted beams is dis-satisfying though. I would have tried to just take the transmitted beam and interfere them via short path lengths and no lenses. The cavity eigenmodes of each cavity should already be matched to ~1%. If the path length from the cavities to the BS can be made equal, the overlap should also be good. In general, it is hard to make a low phase noise setup with a long path length.

To ensure low backscatter from the transmission RFPDs back into the cavity, the beat signal PD should be placed slightly ahead of the focus of the final lens.

And since the Z=40 kOhms for the 1811, the light power from each cavity should be made ~200 mV / 40000 ~ 5 uW. I know its tiny, but otherwise the 1811 is not going to be very linear. And the RF signal going into a level 13 mixer ought to be less than 300 mVrms.

The better option is to use a RFPD with only ~1 kOhm of RF transimpedance...

  289   Wed Aug 18 00:07:37 2010 FrankDailyProgress noise level of beat signal

i tried a couple of things today to figure out what determines the low frequency noise level of the beat measurement.
I did the following steps, nevertheless the noise level didn't change:

  • re-aligned everything starting with the PMC
  • coupling into the cavities is >90%
  • matched power levels in both paths to match power at RF photodetector
  • re-positioned the RF photodiodes for locking the caviries
  • added beam dumps for beams reflected from RF photodiodes
  • optimized overlap of both beams on RF photodiode (beat)
  • changed mixer from 7dbm model to 13dbm model
  • removed cable going to DAQ from DC output of RCAV RF photodiode as it is causing ~36MHz oscillation of RF PD (even with no light)
  • changed gains for both loops, ACAV and RCAV
  • added first QPD at pick-off right in front of periscope into chamber for RCAV to check pointing

things to be checked tomorrow:

  • power fluctuations of beams into both cavities
  • power fluctuations of transmitted beam
  • pointing of both beams (should also show up in power noise spectrum in transmission)
  • measure TF of all loops with current setting for noise estimation

other things to change:

  • move temp readout to VME based stuff
  • rename VCO input monitor signal channel name
  • remove PSL RT stuff from fb0 to see if networking problems are caused by that
  • replace current cables for temp readout due to loose connection somewhere on table
  • check ACAV RF photodiode
  • change PID controller variable names in order to add other software loops like feedback from VCO input signal to RCAV temp setpoint. Right now we have already 4 software loops:
  1. feedback to laser for RCAV using FAST actuator signal
  2. temp ctrl for RCAV
  3. temp ctrl for ACAV
  4. feedback to laser temp for ACAV using VCO input signal (used to track ACAV temp tuning to match RCAV resonance)

 

  288   Tue Aug 17 18:54:19 2010 FrankNotes list of current problems and things to do

 - personal notes -

  • tiny, periodic spikes in acav RF PD
  • ~36MHz oscillation (without light too) at RCAV RF pd dc output
  • PMC transPD signal into DAQ broken, changes periodic from real value to some random number. However real signal at input of cross-connect panel is constant (on scope & multimeter)
  • kicking the table introduces 5mK jumps in temp signals (random channels)

to-do-list (other stuff)

  • switch temp readout to VME based system (3123, channel 9-16, J5 and remaining channels from J4)
  • improve PID parameters (so I and D so far)
  • exchange linux workstation (freezes from time to time)
  • update boot values for all channels
  • change channel name for VCO input signal
  • change channel names for all PID controllers
  • measure TF for all loops
  • ...
  287   Tue Aug 17 15:57:11 2010 FrankNotesBEATpower on beat PD

- personal notes -

New Focus PD:

power from ACAV: 0.958mW
power from RCAV: 0.967mW

DC-OUT: 1.94V
AC-OUT: 1.27Vpp @160MHz in 50R

  286   Tue Aug 17 14:52:57 2010 FrankPhotos beat setup panorama picture incl beams

beat_panorama_small_with_beams3.jpg

  285   Tue Aug 17 14:35:34 2010 FrankSummaryRefCavACAV re-aligned

re-aligned the ACAV once more, now about 91% are coupled into the cavity

  284   Mon Aug 16 19:40:52 2010 taraNotesNoiseBudgetMATLAB simulation 1D

After my fail attempt on simulating those thin layers of coatings on COMSOL, I turned back to MATLAB.

The model is based on 1-D heat equation. I haven't add the effect of coating yet. This will be done soon after.

I assume no radiative loss on any surface, only the effect from AC input. The boundary condition is sinusoidal on the

reflecting surface of the mirror, and no heat flow at the back of the mirror (dT/dz = 0).

Attachment 1: contourplot.png
contourplot.png
Attachment 2: T_Z.png
T_Z.png
Attachment 3: heatflow.m
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%% Script to solve the heat equation substrate with
%%% sinusoidal T at the surface
%%% 2010 Aug 16,
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

%%% Initialize variables
dz = 1e-7; %each depth step is 0.1 um
Nz = 400; % Choose the number of depth steps (should go to at least 100 um)
Nt = 5000; % Choose the number of time steps
... 48 more lines ...
  283   Mon Aug 16 16:36:35 2010 FrankSummaryLaserlaser, PMC, refcav re-characterization and recent changes

laser & PMC:

re-characterized the laser and PMC

new laser temp setpoint : 44.1 C  - almost centered between mode hops,
also PMC resonance close to that with centered PZT offset

  • PMC PZT OFFSET (PMC_RAMP) = -2.7V = 145V @ PZT (300V max)
  • PMC PZT tuning: 2.24V / FSR (714MHz)  for EXT DC INPUT
  • PMC TRANSPD close to limit: changed gain, now 3.32V=24.6mW  (7.41mW/V)

current values for TEM00:
RCAV resonant at 0.0182V
ACAV resonant at 0.0566V

ACAV:

  • re-aligned and re-modematched ACAV as old mode matching was only 47%
  • now DC voltage in reflection is 359mV if not locked and 42mV if locked. -> now 88% coupled into cavity
  • position of lenses very sensitive -> changing position a millimeter makes huge difference but no space for different mode matching
  • changed position of RF photodiode a little bit
  • beam size on PD now a little bit smaller
  • beam dump added for reflected beam from RF photodiode
  • changed setpt temp to 36.5 C, since the box changed the difference between both cavities. So the actual temp read and stabilized is not the (only) one seen by the cavity. The gradient seems to be important too.
  282   Fri Aug 13 20:23:33 2010 TaraNotesNoiseBudget 

I tried to separately mesh the geometry.

Only the small part (1/2 radius, 1/3 thickness) is extremely fine mesh, the rest is regular mesh, see fig1.

The result looks similar, with shorter time for calculation.

Also, the contact area between the mirror and the spacer

is modeled as a constant temperature (@ 35 C)boundary condition.

 The temperature profile across the surface,

nd the temperature difference along the depth due to RIN is plotted below.

The area under the curve for T,Z is 6.1e-9 [meter.Kelvin]

 Expansion coeff for SIO2 is 5.1e-7. Thus, dL = 3.1e-15m.

Use df/f = dL/L, L =0.2035m, we get df = 4.5 Hz.

 

As for thin coatings, I tried swept mesh, which creates prisms of quadrilateral surface.

However, when I ran the simulation, there are warnings about

ill-precondition from the size of the coatings.

After a talk with Greg's SURF student, he wrote a code similar to what I'm doing.

I'll try MATLAB code this time.

Attachment 1: mesh.png
mesh.png
Attachment 2: contact.png
contact.png
Attachment 3: TR.png
TR.png
Attachment 4: dT_Z.png
dT_Z.png
  281   Fri Aug 13 17:03:13 2010 MeganPhotos The Foam Box

The new foam box has a hat! Or lid... Whichever you prefer.

100_1070.JPG

  280   Fri Aug 13 15:47:36 2010 MeganSummaryNoiseBudgetNoise Budget

Quote:

Quote:

Sorry, here's a prettier graph! I think the strange bumps in the low frequencies is from very few samples in that region and the interpolation matlab is trying to do. Once I have more data to compare it to, I'll try to figure out how to get rid of them.

 It looks a little like the FFT was done wrong here. How about uploading some more details, e.g. what is the channel names used? Is this in the DAQ or just using the SR785? How about a diagram showing what the transmission PD is, what the range of the Marconi is, how the signal goes to the DAQ, etc, etc, etc.....

Don't we need a whitening filter to put the PLL control signal into the DAQ? Or I guess we can use a DC channel and an AC coupled, high gain channel.

 This was directly from the SR785, where I saved the data using the GPIB scripts. This is the data with the 1MHz input range of the Marconi with the New Focus photodiode signal. The two signals go into a mixer, through a 1.9MHz LPF, then to the Stanford SR560 with a gain of 2000 and a break frequency of 0.03Hz on the low pass filter (the values necessary for it to actually lock). This feeds back to the input of the Marconi. We looked at data from the error point and the feedback loop with an input range of 100kHz to compare, calibrating the error point by pi/Vp-p*UGF (compensated for the gain by multiplying by 1/f, but then also multiplied by f to give the frequency noise) where Vp-p was 836mV and UGF was 16.2kHz as measured by taking a transfer function with the SR785. The feedback loop was calibrated by multiplying by 71kHz/V, the measured range of the Marconi (measured by applying 0V and 1V and finding the frequency difference). The 1MHz data was from the feedback loop and calibrated by multiplying by 715kHz/V (measured range at 1MHz setting). All the curves showed the same features and were at roughly the same level (within a factor of 2). I chose to plot the 1MHz data because it extended to lower frequencies (possibly a poor choice). We could not get lock with 100kHz input range for long enough to measure the lower frequencies because the cavities were drifting too much in temperature. We looked at the channels for temperature, and it looks like a 1K step in room temperature was causing the cavities to drift quite a bit. So our goal is to restabilize the cavities and take another measurement to verify the ones we already took.

*EDIT* Frank pointed out a concern with the lower frequencies. I attached a graph with the actual data. The strange spikes in low frequencies is from the interpolation of data that isn't real to begin with because of the measurement process. I'm working to figure out how to fix the interpolation and will hopefully have a graph of something less confusing soon. Also, the flat line at higher frequencies is electronic noise, we believe from the demodulation setup, because the line is always present, but higher with higher input ranges.

*EDIT2* The noise budget is back! I just deleted the low frequencies that aren't accurate and the high frequencies that weren't measured because I haven't figured out how to accurately extrapolate what we have.

Attachment 1: BeatSignal.png
BeatSignal.png
Attachment 2: NoiseBudget.png
NoiseBudget.png
Attachment 3: NoiseBudget.fig
  279   Fri Aug 13 15:10:33 2010 ranaSummaryNoiseBudgetNoise Budget

Quote:

Sorry, here's a prettier graph! I think the strange bumps in the low frequencies is from very few samples in that region and the interpolation matlab is trying to do. Once I have more data to compare it to, I'll try to figure out how to get rid of them.

 It looks a little like the FFT was done wrong here. How about uploading some more details, e.g. what is the channel names used? Is this in the DAQ or just using the SR785? How about a diagram showing what the transmission PD is, what the range of the Marconi is, how the signal goes to the DAQ, etc, etc, etc.....

Don't we need a whitening filter to put the PLL control signal into the DAQ? Or I guess we can use a DC channel and an AC coupled, high gain channel.

  278   Fri Aug 13 11:06:10 2010 MeganDailyProgress New Layer of Foam for the Cavities

We were going to work on building a metal box around the cavities, but they didn't have enough sheet metal and would have to order more, so Frank is working on a design for the box and will order the metal.

Instead, we got free foam! Someone ordered too much, so we took a sheet to build a new box around the cavities to see if it would help with stabilization.

After a while spent cutting the foam, we built a box!

100_1054.JPG

100_1058.JPG

100_1063.JPG

And it is very happy to stabilize the cavities. We placed some foam bits on top to make a makeshift lid, and will use another foam sheet today to cut a real lid that fits on the box.

  277   Fri Aug 13 01:14:53 2010 FrankNotesNoiseBudgetDC simulation

graph?

Quote:

I used COMSOL to simulate temperature change caused by 10^-4 fluctuation from the main DC power.

That is the temperature change in the substrate due to power change from Pin to Pin + Pin*10^-4.  Pin is 10mW.

This should give us the upper limit of the expected frequency noise.

The area under the plot of Temperature vs Depth is 1.9 * 10^-8 [meter*Kelvin].

 

thermal coefficient for substrate is 0.51 *10^-6

thus dL is (0.51*10^-6) * (1.9*10^-8) ~ 10^-14.

Use dL/L  = df/f ;  where L is the cavity length = 0.2035m,  f = c/ lambda.

df = 15 Hz. (effect from thermal expansion 1 mirror in the cavity only)

 

  276   Thu Aug 12 17:34:37 2010 MeganSummaryNoiseBudgetNoise Budget

Sorry, here's a prettier graph! I think the strange bumps in the low frequencies is from very few samples in that region and the interpolation matlab is trying to do. Once I have more data to compare it to, I'll try to figure out how to get rid of them.

Attachment 1: NoiseBudget.fig
Attachment 2: NoiseBudget.png
NoiseBudget.png
  275   Thu Aug 12 16:27:09 2010 taraNotesNoiseBudgetDC simulation

I used COMSOL to simulate temperature change caused by 10^-4 fluctuation from the main DC power.

That is the temperature change in the substrate due to power change from Pin to Pin + Pin*10^-4.  Pin is 10mW.

This should give us the upper limit of the expected frequency noise.

The area under the plot of Temperature vs Depth is 1.9 * 10^-8 [meter*Kelvin].

 

thermal coefficient for substrate is 0.51 *10^-6

thus dL is (0.51*10^-6) * (1.9*10^-8) ~ 10^-14.

Use dL/L  = df/f ;  where L is the cavity length = 0.2035m,  f = c/ lambda.

df = 15 Hz. (effect from thermal expansion 1 mirror in the cavity only)

  274   Thu Aug 12 11:09:56 2010 MeganNotesNoiseBudgetBeat Added to Noise Budget

As Frank said, we tested the GPIB yesterday. I took one set of data that we got and put it in the noise budget to see how it compares. We will also hopefully take more measurements today with newly adjusted stuff so I can put that measurement in instead if it differs.

Attachment 1: psl_refcav_sio2_300K.png
psl_refcav_sio2_300K.png
  273   Thu Aug 12 00:15:59 2010 Frank, MeganSummaryComputersnetgpib scripts working

Megan tested the gpib scripts today and they are working except that the plotting of the data doesn't work as some packages are still missing.
Unfortunately i couldn't install them with yum so i will do it later.
Anyway, taking data without the plot function is working ...

  272   Thu Aug 12 00:10:44 2010 FrankSummaryRefCavboth cavities locked ~24h

this afternoon both cavities were locked since 24h. The initial problem was to find the right temperature for one of the cavities if we keep the other one fixed at 35 degrees. So now the reference cavity is held at 35 degrees, (because the time constant it needs to settle is much higher), while the analyzer cavity is at 35.3 degrees. With those setting the VCO input signal is almost at zero, having about +/-10MHz of tuning range.

  271   Wed Aug 11 23:57:10 2010 Frank, PeterDailyProgressFSSFSS loop fixed

Peter and i solved the problem with the fss loop today, but here the long story:

The problem was that the loop was much more stable without the PC connected, so only the PZT of the laser was used as an actuator so far. Already some time ago i thought it might be the wrong sign for the PC and so i tried to change the sign of the loop by changing the two jumpers at the output going to the fast actuator and changing the sign of the error signal, but i never got it to lock in that configuration. So i though the other, previous configuration must be the right one, as it is locking. Later Tara had the same problems. The problem was that if the PC was connected the stability didn't increase and it even seemed to be more unstable using the PC in addition.

So in fact, the sign we used so far was the wrong one. But the problem was that changing the sign of the fast actuator and then flipping the error signal sign didn't work because the other (right) jumper settings don't work ! There is no feedback to the PZT of you change the two jumpers to the other orientation !! That's why it never worked with the right jumper positions.

As we figured out that this might be the problem i've build a BNC adapter changing the sign at the input of the laser and here we go, it's working now  So something is wrong with this board . I will figure that out later what exactly is broken on that board.

will measure the ugf of the fss loop tomorrow...

  270   Tue Aug 10 22:57:25 2010 FrankDailyProgressRefCavstrange channel behavior

we have probably some minor network problems which causes some channels not beeing available at all times even on the same computer.

example:
if you open the Striptool and Probe at the same time and access the same channel, one of the tools can see the channel, the other doesn't.
Or one tools sees a different value, e.g. if i change the setpoint of the temp loop, probe still sees the old value, but the Striptool already the new one.
Sometimes it helps if you close the application and after restarting it everything is fine. Same for the framebuilder. Sometimes some of the channels get lost.
I don't know where it comes from. It happens randomly to any of the channels. It's not the traffic on the network, but could be some routing problem.

  269   Tue Aug 10 21:40:42 2010 Megan and FrankDailyProgress Both Cavities Locked!

We got both cavities locked today!

They stayed locked for at least an hour and were still locked when we left. Hopefully they'll still be locked tomorrow morning!

We aligned both beams into the photodiode (they were already pre-aligned) and maximized the alignment to get the largest signal out of the photodiode. We locked the signal from the photodiode with the IFR/Marconi 2023B signal generator and took some measurements of noise and transfer functions to determine UGFs, which I will post graphs of when I put them together. The plan is to let the temperatures really settle down (they were still settling a bit when we left) and take a good measurement of the noise and measure the transfer functions to determine UGF tomorrow. Then we will pull everything apart (not really) to replace the two cables that have extensions on them to see if something in the connection can be fixed to give better stabilization.

*I added a graph of the noise we measured yesterday at 3 different times*

Attachment 1: BeatSignal.png
BeatSignal.png
  268   Tue Aug 10 19:35:13 2010 taraNotesNoiseBudgetfluctuation in power, COMSOL model

I use time dependent option to simulate temperature change inside the mirror. The power input is DC + fluctuation.

Assuming RIN = 10^-4 @ 10 Hz, no coatings layer.

The plot below shows the absolute value of temperature fluctuation |(T - <T>)| on y-axis, and depth from the surface on x-axis.

Different lines (data 1 to 11) are different times in the substrate, from 20.0 to 20.1 sec with 0.01 time step.

It seems that the thermal length is quite large ~0.26 mm compared to what we got from half infinite model (~5 um).

 

dT from 1-D model is ~0.32 uK, while the result from COMSOL is ~1.8 uK.

If we use dT = 1.8 uK the frequency noise at 10 Hz will be (a factor of 6 larger) ~0.009 Hz/ rtHz. Still too low to see.

Attachment 1: FIT_thermal_length.png
FIT_thermal_length.png
Attachment 2: heat3t.mph
  267   Tue Aug 10 01:10:35 2010 FrankSummaryRefCavupdated schematics of the setup

entire setup (simplified version):

refcav-setup_v1.png

 

beat signal generation in more detail:

refcav-beat_v2.png

  266   Mon Aug 9 19:57:29 2010 taraNotesNoiseBudget DC simulation by COMSOL 4.0

I use COMSOL to simulate the mirror of RefCav.

I tried only DC heating from the beam. The mirror has finite size, absorbed power is radiated outside.

Next step will be using DC power + small fluctuation.

 The information about dT will be used to calculate phase shift due to thermoelastic and thermorefractive noise.

 

 

note about heat3.mph file

 COMSOL 4.0

MIRROR model for PSL
This is a note about file heat3.mph

1) Global definition
     1.1) parameters: most of them are self explanatory. Some that might need
                     clarification are
             a)  p0 is the total power absorbed on the surface
                  which is Power input x absorption x finess/pi 

                  = 10mW x 5ppm x 10^4/pi ~ 1.6 x10^-4 watt

             b) I0, (p0/2pi w^2) Intensity factor of the Gaussian beam, when integrated the
                   intensity over  the area, the total power will be p0

              c) Qin, gaussian beam profile on the mirror

2) Model 1

   2.1) Geometry 1
           a) a substrate with thickness, radius as specified in (1.1)
           b, and so on) multiple coating layers*
              *note: I tried both 3 doublets with 1 cap and effective 5 um coating, but when I mesh the geometry, it's out of memory.
              So I disabled all coatings, and use only substrate. Frank suggested I can reduce the
              mirror's size if the temperature across the surface drops fast enough. I haven't tried that
              *effective coating is a single layer, 5 um thick. Thermal expansion coeff and heat capacity is
               calculated from multiple layers of Ta2O5 and SiO2.
 
          c) Form Union-> choose Form a union then, click "Build All"  

  2.2) Materials
        a) SiO2 (fused quartz): from COMSOL library
             surface emissivity is 0.75 (taken from  <http://www.optotherm.com/emiss-table.htm> and

<http://www.google.com/url?sa=t&source=web&cd=2&ved=0CCIQFjAB&url=http%3A%2F%2Fwww.holanengineering.com%2Fsitebuildercontent%2Fsitebuilderfiles%2Femissivity_1.pdf&ei=JGBgTKr9IIL2tgOPj-ipCw&usg=AFQjCNHFASfDeyG6yvq-FNXuEUShtmH_2Q>

    b) Ta2O5, all parameters are from Evans etal paper.
            Note that the value from Evans' paper,
           the specific heat capacity for Ta2O5 is  2.1 X10^9 [Joule/m^3 kelvin]
             in comsol I defined
                specific heat capacity to be 2.1 x10^9[Joule/kg Kelvin] and,
                 mass density to be     1.0   [kg/ m^3] , so heat capacity per volume is the same.
     c) effective coating's properties from multiple layers of SiO2 and Ta2O5
   thermal conductivity
 
2.3) Heat Transfer(ht)
         a) heat transfer in solid: all domains
         b) thermal insulation : only coatings' edge
        c) initial values: all domains are set at 310 K
         d) boundary heat source: at the topmost surface, Qin is defined in (1.1) (gaussian beam)
         e) surface to ambient radiation: all surfaces except coatings' side
              *for surface emissivity value, I have to choose User difined to be 0.75. I tried "from material"
                but it didn't work even though I add the surface emissivity in material property already, see (2.2)
2.4) Mesh
       - choose free tetrahedral, and specify size as you want, click build all.

3) study
    a) right click on study, select compute. Sit back, relax and enjoy the result.

4) Results
    a) 3d plot group1 shows a 3D picture
    b) 3dplot Group2 shows a slice
    c) 1D temp center axis shows a temperature plot along the center axis

Attachment 1: Pic1.png
Pic1.png
Attachment 2: temp_across.png
temp_across.png
Attachment 3: Temp_depth.png
Temp_depth.png
Attachment 4: heat3.mph
  265   Mon Aug 9 17:34:43 2010 FrankSummaryRefCavtemperature jumps

the temperature in some of the cavity channels jumps about 5mK if you kick the table hard enough, so there is some problem with the d-sub connections on the table connecting the sensor readout box with the DAQ.

 tempsteps.png

  264   Mon Aug 9 12:55:40 2010 FrankNotesRefCavslow actuator values for both cavities

both cavities @35 degrees +/- 5mK

 

slow actuator values

RCAV : 0.1890
ACAV : 0.1993

(.1890 V - .1993V) * 1180 MHz/V =  12 MHz difference

  • temperature of one cavity has to be tuned about ~12MHz with 6mK/MHz, so ~70mK
  • loosing lock @~+3V input to the VCO which should go up to at least 5V

anyway, the tuning range is only 10MHz in that direction, so can't lock both at the same time right now

frequency see here: VCO tuning

  263   Mon Aug 9 00:21:39 2010 FrankSummaryRefCavThe low UGF on Frank's temperature servo

What about the idea of putting both cavities in one chamber? Then we don't have to worry about the temp stability anymore.

Quote:

Dear Frank,

I think basically you're right about the foam and the RC can's temperature servo. Although the impulse response of the can to heat is fast, the cool down time must be slow as its dominated by the conduction through the foam. Since this makes the servo asymmetric, it cannot have a high gain at the 24-hour period, unlike the system at the 40m.

So, what's the answer? We need to be able to keep the 24-hour period to be less than 5 MHz in the relative frequency shift. Ignoring the effects from the coating we ought to get ~150 MHz / K. At the 40m, we actually see more like 110 MHz / K at DC.

Assuming that you will always have 3 deg p-p fluctuations down there, we need a gain of 60 at 24 hours. This means that the UGF should be ~24 hours / 60 = 20 minutes. So that sets the maximum amount of foam that we are allowed to have on the can.

That ought to allow us to make the measurement for the first phase (the one with the double AOM setup).


For the second stage, we ought to use your Aluminum box idea. Put a 2 mm thick Al box completely enclosing the RC can leaving 4-5 inches of space. Layer the outside of the Al box with a 2-3 cm thick sheet of foam to keep the heating power low. In that setup, we ought to be able to go wild with the RC foam since the DC control will all be done with the faster Aluminum heater.

Rana

  262   Sat Aug 7 22:01:45 2010 ranaNotesNoiseBudgetRIN induced Photo Thermal Noise calculation

Quote:

I used the result from [1] to calculate frequency noise from thermoelastic and thermorefractive due to RIN.

at 10 Hz, it is about 1.4mHz/rtHz. RIN at 10 Hz used here is 10^-4.

This calculation does not include power fluctuation from RFAM/ pointing stability.

The result will be compared with a simulation from COMSOL (in progress.)

[1] Evans et al, PHYSICAL REVIEW D 78, 102003 (2008)

 FYI:    Upper Limit meaured at the 40m   4 months ago.

  261   Fri Aug 6 23:41:36 2010 ranaSummaryRefCavThe low UGF on Frank's temperature servo

Dear Frank,

I think basically you're right about the foam and the RC can's temperature servo. Although the impulse response of the can to heat is fast, the cool down time must be slow as its dominated by the conduction through the foam. Since this makes the servo asymmetric, it cannot have a high gain at the 24-hour period, unlike the system at the 40m.

So, what's the answer? We need to be able to keep the 24-hour period to be less than 5 MHz in the relative frequency shift. Ignoring the effects from the coating we ought to get ~150 MHz / K. At the 40m, we actually see more like 110 MHz / K at DC.

Assuming that you will always have 3 deg p-p fluctuations down there, we need a gain of 60 at 24 hours. This means that the UGF should be ~24 hours / 60 = 20 minutes. So that sets the maximum amount of foam that we are allowed to have on the can.

That ought to allow us to make the measurement for the first phase (the one with the double AOM setup).


For the second stage, we ought to use your Aluminum box idea. Put a 2 mm thick Al box completely enclosing the RC can leaving 4-5 inches of space. Layer the outside of the Al box with a 2-3 cm thick sheet of foam to keep the heating power low. In that setup, we ought to be able to go wild with the RC foam since the DC control will all be done with the faster Aluminum heater.

Rana

  260   Fri Aug 6 22:24:46 2010 FrankNotesComputersmissing file for medm screens

- personal notes -

medm error message: loadGIF: Cannot open file:  /cvs/cds/caltech/scripts/utilities/nova_logo.gif

  259   Fri Aug 6 14:36:24 2010 Megan, FrankNotesRefCavcurrent slow actuator values

- personal notes -

slow actuator values

RCAV : 0.1924
ACAV : 0.2043

(.1924 V - .2043V) * 1180 MHz/V =  14 MHz difference

  258   Fri Aug 6 11:51:47 2010 MeganDailyProgress VCO Measurements

Yesterday we set up channels to record the noise of the locked signal generators. After a while playing with channels and filters and everything, the results are that the missing factor might be somewhere in the calibration of the data in 40m because this data fits where we would expect it to, slightly above the electronic noise converted to phase noise at low frequencies and much higher at high frequencies.

I tried to check the calibration in 40m this morning, but Jenna had the rubidium clocks hooked up and was collecting data with that, but I might be able to go back this afternoon and get something to see if it's in the calibration of rad/count.

Attachment 1: NoiseComparison.png
NoiseComparison.png
  257   Fri Aug 6 00:43:09 2010 FrankSummaryComputerswifi bridges configured

as Mott can't finish the installation of the wifi bridges for the gpib-to-ethernet adapters a configured 4 devices today.
Username and password are clearly labeled on each device, they are the usual ones for administrator rights.

The 4 new devices have the following IP adresses:

10.0.1.6
10.0.1.7
10.0.1.8
10.0.1.9

i will finish configuring the gpib-to-ethernet adapters on Friday, one for each instrument in the labs.

i will update the network diagram as soon i decided which ip addresses the will get

  256   Thu Aug 5 11:33:20 2010 MeganDailyProgress BluePhase 1000

The cavities were still too far off to lock simultaneously, but they should hopefully be better today.

Instead, we played with the Wenzel BluePhase 1000 phase noise test system.

P1550991.JPG

Following the directions in the manual, we locked an IFR/Marconi 2023A to an IFR/Marconi 2023B, calibrated according to the manual, and got a curve very similar to what we had before . This means it's probably working correctly!

So then, just to be adventurous, we decided to connect the VCO in place of the 2023B to measure its phase noise. Apparently that was too adventurous for the BluePhase 1000 . We couldn't lock it (with the feedback loop going to the 2023A) with any input range below roughly 100 kHz and the output (after calibration as per the manual) was a flat line. After readjusting cables and reconnecting cables and finally reverting back to the 2023A vs 2023B to make sure the machine still worked, we decided the unity gain frequency was probably pretty high and the manual calibration did not take that into account. So we used a swept sine measurement to find the transfer function of the system with the VCO connected and locked. The UGF was around 5.3 kHz with the 100 kHz input range . This means that the calibration doesn't account for everything when the UGF is high . But it also means we may have found the problem with our data when the VCO is connected ! So I have to take the data we have, apply a zero at 5.3 kHz, and see if that gets it to line up correctly.

Meanwhile, we installed a program on the computer that, when connected to the BluePhase 1000, can control all the knobs and buttons and locking remotely. And we discovered you can do more on the computer than with the switches on the front! Like change the capacitor value.

So the summary of yesterday's activities is: don't ever completely trust the calibration the manufacturers tell you to use. They might not be taking something (UGF) into account.

And the calibration as per the manufacturers:

1) Adjust the offset until exactly 1 period is displayed on the oscilloscope

2) Divide the time divisions on the oscilloscope to 1/100th the original (this gives 0.02Pi radians)

3) Measure the voltage difference across the two ends of the line

4) Calculate your slope! (gives V/rad)

5) Noise [dBc/Hz] = [PSD]-[20log(slope)]-[amp gain]-[correction for SSB measurement]

Attachment 1: GraphCompare.png
GraphCompare.png
  255   Thu Aug 5 11:17:18 2010 taraNotesNoiseBudgetRIN induced Photo Thermal Noise calculation

I used the result from [1] to calculate frequency noise from thermoelastic and thermorefractive due to RIN.

at 10 Hz, it is about 1.4mHz/rtHz. RIN at 10 Hz used here is 10^-4.

This calculation does not include power fluctuation from RFAM/ pointing stability.

The result will be compared with a simulation from COMSOL (in progress.)

[1] Evans et al, PHYSICAL REVIEW D 78, 102003 (2008)

Attachment 1: RINnoise.m
c=3e8;          %speed of light
lambda = 1064e-9;

Finesse = 1e4;
L=0.2035;       %cavity length

%adjustable parameters
f=10;           %Let's look at 10 Hz
Pin= 10e-3;     % power input 10 mW
RIN= 1e-4;      % RIN level at 10 Hz
... 40 more lines ...
  254   Wed Aug 4 16:21:01 2010 FrankNotesRefCavslow actuator values for both cavities

 - personal notes -

slow actuator values

RCAV : 0.1915ACAV : 0.2050

(.1915 V - .2050 V) * 1180 MHz/V = -15.93 MHz difference

  253   Wed Aug 4 10:56:04 2010 MeganDailyProgress Optics Set Up!

All of the optics were set up in the path of the transmitted light from both cavities. Each path needed both a quarter wave plate and a half wave plate to be able to get the s polarization needed by the beam splitter. This solved problems we had earlier with the 50/50 beam splitter doing more like 25/75!

For mode matching, 2 lenses were placed in the path of the RCAV, one with a focal length of 175mm at a distance of roughly 7.5 inches from the first mirror and the other with a focal length of 200mm roughly 5.5 inches before the second mirror. This gave a waist of 224 µm 7.5 inches after the beam splitter.

The ACAV got only 1 lens, with a focal length of 200mm, 2.5 inches before the second mirror, with a waist of 208 µm at the same distance (7.5 inches from the beam splitter).

The first photodetector was placed with a mirror (for better alignment) above the beam splitter, and the other was placed after 2 mirrors and another lens in the other path. The second photodetector has a smaller detector area, so we ideally would like the waist at the detector to be under 100 µm. To make the beam smaller, we used another lens with a focal length of 30mm roughly 2 inches after the second mirror. This still leaves the beam a little larger than we'd like (it's very sensitive to alignment), but it will be fine for the moment.

The two cavities have still not been intentionally locked simultaneously. Yesterday, they differed by roughly 17 MHz (if I remember correctly) and so they were being heated overnight so their temperature can stabilize and we can see if they are closer today.

  252   Tue Aug 3 14:41:59 2010 FrankNotesPurchaseslenses to replace in kit

KPX088  75.6mm

KPX100  150mm

KPX103  175mm

KPX106  200mm

KPX109  250mm

  251   Tue Aug 3 11:22:04 2010 MeganDailyProgress PSL Table Updates

Both the analyzer cavity and reference cavity are able to be locked, and were simultaneously locked for a while yesterday . However, the temperature control loop is not working because the computer is not seeing the temperature data, so it didn't stay locked . But now we know it's possible !

The mirrors, beam splitter, and photodetectors are set up and aligned to measure the beat of the two signals and I am currently working on mode matching to get the beams the same size at the photodetector.

FSS_RCTRANSPD and ACAV_TRANSPD have been calibrated:

FSS_RCTRANSPD = 1.98 mW/V

ACAV_TRANSPD = 2.18 mW/V

  250   Mon Aug 2 16:23:59 2010 FrankNotesDAQconnectins for PMC, RCAv and ACAV to VME-based system

- personal notes -

current cross-connect connections

 

------------  PMC --------------

C3:PSL-PMC_RFPDDC:
block TB4
1 - LO
2 - HI
connected to J3-3113A, 1/2 (CH32)

-------------

C3:PSL-PMC_PMCERR:
block TB4
3 - LO
4 - HI
connected to J3-3113A,  5/6 (CH34)

-------------

C3:PSL-PMC_PMCTRANSPD:
connected to J4-3123  5/17 (CH2)
5 - LO  /  17 - HI

 

------------  RCAV --------------

C3:PSL-FSS_RFPDDC:
block TB2
1 - LO
2 - HI
connected to J2-3113A  43/44 (CH21)

-------------

C3:PSL-FSS_RCTRANSPD:
connected to J4-3123  2/14 (CH0)
2 - LO  /  14 - HI

-------------

C3:PSL-FSS_MIXERM
block TB2
5 - LO
6 - HI
connected to J2-3113A  53/54 (CH26)

 

------------  ACAV --------------

C3:PSL-ACAV_RFPDDC:
block TB2
3 - LO
4 - HI
connected to J2-3113A  47/48 (CH23)

-------------

C3:PSL-ACAV_RCTRANSPD:
connected to J4-3123  16/3 (CH1)
16 - LO  /  3 - HI

 

------------  TEST --------------

C3:PSL-TEST_SHORT:
connected to J4-3123  25/12 (CH7)
25 - LO  /  12 - HI

 

  249   Mon Aug 2 12:16:04 2010 MeganNotes VCO Model on SVN

The model I made of the VCO is now on the svn, in the filter folder in liso, listed as VCODriverNoise.fil

  248   Fri Jul 30 13:23:29 2010 FrankNotesDAQVMIC 3123 connections (J4 and J5) (16bit adc)

- personal notes -

list of current connections to the VMIC 3123 card in the PSL rack:

J-4:

LO/HI

2/14 : FSS_RCTRANSPD

16/3 : ACAV_RCTRANSPD

5/17 : PMC_PMCTRANSPD

19/6 :

8/20 :

22/9 :

11/23 :

25/12 :

J-5: no connections

 25/12 : "TEST" , test input, shorted

 

list of old connections to the VMIC 3123 card in the PSL rack according to document D980535-C-C:

J-4:

2/14 : FSS_MIXERM

3/16 : FSS_SLOWM

5/17 : PMC_PMCTRANSPD

6/19 : FSS_MINCOMEAS

8/20 : ISS_ISERR

9/22 : PMC_PMCERR

12/25 : FSS_RMTEMP

 

J-5:

2/14 : FSS_RCTEMP

3/16 : FSS_RCTRANSPD

  247   Fri Jul 30 13:17:05 2010 FrankSummaryComputersRT code built using wrong user
i deleted the PSL compiled stuff and rebuilt it with the controls user. Someone wants to do this for the ATF
model as well...

<p>
<table width="98%" cellspacing="1" align="center" style="border: 1px solid rgb(72, 96, 144);">
    <tbody>
        <tr>
            <td style="background-color: rgb(72, 96, 144); color: white;" cellpadding="3px">Quote:</td>
        </tr>
        <tr>
            <td style="background-color: rgb(255, 255, 176);" cellpadding="10px">
            <p>I noticed that root built the last ATF model and this screwed some things up. I am not sure if
this was also done with the PSL model, and if it might somehow cause network problems.</p>
            </td>
        </tr>
    </tbody>
</table>
</p>
<p>&nbsp;</p>
  246   Thu Jul 29 13:08:40 2010 taraNotesRefCavAlignment is Off

This morning Megan tried to lock RefCav, but the alignment was off. I realigned it, and now it's locked.

C3:PSL-MGAIN  4.4 dB

C3:PSL-FAST 16.6 dB

 

  245   Thu Jul 29 02:55:15 2010 DmassSummaryComputersboth cavities individually locked

I noticed that root built the last ATF model and this screwed some things up. I am not sure if this was also done with the PSL model, and if it might somehow cause network problems.

  244   Thu Jul 29 01:25:42 2010 Frank, MeganDailyProgressRefCavboth cavities individually locked

we locked both cavities individually. Right now we can't lock both at the same time as the temperature stabilization loop isn't working. Something is wrong with the network.
We don't see the channels taken with the realtime system (running on fb0). Something is wrong with the broadcasting of the epics channels.
So we decided to lock them individually for now and start setting up the beat measurement as far as we can. The problem might be solved until we really need both at the same time.
We also add some more channels for the second cavity, RFPD DC out and transmitted light in order to monitor the quality of lock over time.

ELOG V3.1.3-